A stunt man drives a car out of the back of a moving lorry. For the stunt to work, the car must be moving at a velocity of −5.00ms−1 the instant it has left the lorry. The lorry is travelling at a velocity of 25.0ms−1. What speed must the speedometer on the car reach before the car leaves the lorry?
1
Expert's answer
2019-10-18T10:08:09-0400

The speed of the car relative stunt is


"V=25+5=30\frac{m}{s}"